OpenStudy (anonymous):

Someone please explain this to me step by step? verify the identity cos(x+pi/2) = -sinx

OpenStudy (anonymous):

kitty its logical you just should imagine it on the tir circle so when you say pie/2 +x so your in the second quad thus you have a negative cosine but since its pie/2 cos becomes sine soo its -sinx

OpenStudy (anonymous):

How would I write that out in numbers?

OpenStudy (anonymous):

if you ues the addition angle formula you get \[\cos(x+\frac{\pi}{2})=\cos(x)\cos(\frac{\pi}{2})-\sin(x)\sin(\frac{\pi}{2})\]

OpenStudy (anonymous):

two of those are numbers you good from there?

OpenStudy (anonymous):

From here which would we factor out?

OpenStudy (anonymous):

comeee on mannn where is your logic why going through all this

OpenStudy (anonymous):

It's our last question and we are bad with math =\